OpenStudy (anonymous):

An infant acetaminophen suspension contains 80mg/0.80mL. The recommended dose is 15mg/kg body weight. How many mL of this suspension should be given to an infant weighing 11 lbs?

OpenStudy (abmon98):

acetaminophen suspension contains 80mg/0.80mL. Convert pound to kilogram, 1 lbs is equal to 0.45359237 kilogram. 1 1bs:0.4535 kilograms 11 lbs: x kilograms solve for x kilograms using cross multiplication x=11*0.4535=4.9895 kg 15 mg/kg= y mg/4.9895 y (mg)=4.9895*1/15 80 mg/0.80 ml= 0.3326mg/ z ml

OpenStudy (anonymous):

Okay, thank you. That's very helpful!

OpenStudy (abmon98):

good to hear that you are welcome :)